OpenStudy (anonymous):

a contractor finds it takes sam tockett 9 hours to construct a wall of a certain size. It takes Jim Lowell 8 hours to contsruct the same wall. How long would it take if they worked together

OpenStudy (anonymous):

Can anyone please help????

OpenStudy (apoorvk):

Let the wall construction be a certain \(x\) amount of work. So, if Sam does that \(x\) amount of work in \(9\) hours, then in \(1\) hour, he can do \(\frac x 9\) work. Similarly, Jim does the same \(x\) work in \(8\) hours, then in \(1\) hour, he can do \(\frac x 8\) work. So, together, their rate of working per hour = \(\frac x 8 + \frac x 9\) Now if this^, is their combined rate of constructing the wall per hour, then how many hours would they take they take to complete the entire 'x' work?? The total work 'x', divided by the combined rate!! Solve, and enjoy!